# Legal Minds, Algorithmic Decisions: How LLMs Apply Constitutional Principles in Complex Scenarios

### v1

@@ Abstract
Figure 1: Our experiment consists of three phases. In the first step, a language model — GPT-4 — is given a dataset of legal cases on bioethics in order to extract the arguments made by three legal parties involved. We then prompt the model to state its own position for each case. We embed both sets of arguments, original and generated, with a custom embedding model finetuned to recognize similar legal stances, which allows us to compute the similarity between GPT-4 and the other parties. This measure can be used to make considerations on GPT-4’s inclination towards more progressive or conservative opinions.
@@ 2 Data
2.1 Collecting constitutional rulings
@@
2.2 Using GPT-4 to extract legal arguments
The legal cases described in the previous section, which constitute our dataset of bioethics issues, are complex documents with little internal organization. To carry out our experiment, we need to be able to separate each party’s opinion and arguments, and isolate all different interpretations of the constitutional articles referenced in the case. The list of articles is provided in Appendix D
We use the GPT-4 model from OpenAI (Achiam et al. 2023) to perform this legal reasoning task, specifically the gpt-4-turbo-2024-04-09 version of the model, chosen for its advanced capabilities in handling complex text and context length capacity. The length of the input documents, and the necessity of including a list of constitutional articles in the prompt, required the use of a model with a large context length. The prompt used for the task is presented in Appendix B.1, and a sample analysis generated by GPT-4 can be found in Appendix E. This task tests GPT-4’s ability to summarize text, identify relevant constitutional principles, and outline the arguments presented by different parties.
@@ 2.2 Using GPT-4 to extract legal arguments
Figure 2: Results of the evaluation of GPT-4’s argument extraction task from Section 2.2. The scores, given on a scale from 1 to 5 according to the rubric in Appendix C, show a consistently good performance of the model on the task.
@@ 3 Methodology
3.1 Argument embedding with contrastive learning
@@
Starting from the breakdown of the rulings produced by GPT-4 as described in Section 2.2, we create pairs $(a_{1},a_{2})$ of legal arguments for each of the rulings, such that both arguments refer to the same legal case, but are made by different legal parties. It only makes sense to compare different interpretations of the same principle, so each pair refers to the same constitutional article or group of articles.
@@ 3.2 Measuring GPT-4’s alignment
The model is instructed to structure the overall argument around its choice of relevant constitutional articles, so that the resulting analysis follows the same structure as the dataset created in 2.2. For each ruling, we sample 5 answers to account for the probabilistic nature of the LLM and measure the consistency of GPT-4’s stance over repeated queries. An example of the answer’s format and content can be found in Appendix E.
We then use the embedding model trained in Section 3.1 to embed each of the arguments, and compute the cosine distance between GPT-4’s arguments and the other legal parties’. As previously mentioned, the embedding model was finetuned on the task of representing legal arguments so that opposing arguments will be further apart in space than concordant arguments. Due to this, we assume that this metric can be used to quantify the difference in alignment between different legal interpretations. Furthermore, since all of the text was generated by the same language model using the same data sources, we can mostly rule out the possibility that the embedding model is capturing stylistic differences rather than differences in legal intepretations. Figure 4 shows the results of this measurement, broken down by constitutional article. The next section discusses the results.
@@ 4 Results
As evidenced by our evaluation in Section 2.2, GPT-4 performs more than adequately on the ruling analysis task. The model is capable of summarising and analysing the text with a sufficient degree of completeness and accuracy, with a limited risk of hallucination, but it might require additional background on the relevant jurisprudence to increase the quality of the output. On the other hand, it’s questionable to assert that the model demonstrates a deep awareness of the content. In fact, it is challenging to determine whether the model is actually performing legal reasoning or if it is merely echoing the content it has been trained on. On one hand, the model seems to correctly decode the legal reasoning presented by different parties, accurately discerning the various interpretations of articles mentioned, and grasping the essence of the different positions advocated. It is noteworthy that it exhibits a high level of comprehension even when the cases are written in technical legal language (legalese), which can be complex given the specific domains they cover, requiring a certain degree of expertise. On the other hand, an alternative perspective may contend that the model has merely performed a basic legal task by outlining the main contents of a given text. This may not necessarily indicate a deep understanding of constitutional principles and their nuanced interpretations; it could be seen as simply paraphrasing the information it has been provided.
In any case, as demonstrated in previous research (Guha et al. 2024; Pont et al. 2023), our experiment shows that GPT-4 can be a useful instrument to accomplish basic legal tasks, such as summarising text and outline relevant arguments, under the supervision of legal experts.
@@ Appendix B Prompts
B.1 Extracting legal arguments from a case
@@
B.2 Asking GPT-4 to take a stance in a legal case

### v2

@@ Abstract
Figure 1: Our experiment consists of three phases. In the first step, a language model — GPT-4 — is given a dataset of legal cases on bioethics in order to extract the arguments made by three legal parties involved. We then prompt the model to state its own position for each case. We embed both sets of arguments, original and generated, with a custom embedding model fine-tuned to recognize similar legal stances, which allows us to compute the similarity between GPT-4 and the other parties. This measure can be used to make considerations on GPT-4’s inclination towards more progressive or conservative opinions.
@@ 2 Data
Collecting Constitutional Rulings
@@
Using GPT-4 to Extract Legal Arguments
The legal cases described in the previous section, which constitute our dataset of bioethics issues, are complex documents with little internal organization. To carry out our experiment, we need to be able to separate each party’s opinion and arguments, and isolate all different interpretations of the constitutional articles referenced in the case. The list of articles is provided in Appendix D.
We use the GPT-4 model from OpenAI (Achiam et al. 2023) to perform this legal reasoning task, specifically the gpt-4-turbo-2024-04-09 version of the model, chosen for its advanced capabilities in handling complex text and context length capacity. The length of the input documents, and the necessity of including a list of constitutional articles in the prompt, required the use of a model with a large context length. The prompt used for the task is presented in Appendix B.1, and a sample analysis generated by GPT-4 can be found in Appendix B.2. This task tests GPT-4’s ability to summarize text, identify relevant constitutional principles, and outline the arguments presented by different parties.
@@
Figure 2: Results of the evaluation of GPT-4’s argument extraction task from Section 2. The scores, given on a scale from 1 to 5 according to the rubric in Appendix C, show a consistently good performance of the model on the task.
@@ 3 Methodology
Argument Embedding With Contrastive Learning
@@
Starting from the breakdown of the rulings produced by GPT-4 as described in Section 2, we create pairs $(a_{1},a_{2})$ of legal arguments for each of the rulings, such that both arguments refer to the same legal case, but are made by different legal parties. It only makes sense to compare different interpretations of the same principle, so each pair refers to the same constitutional article or group of articles.
@@
The model is instructed to structure the overall argument around its choice of relevant constitutional articles, so that the resulting analysis follows the same structure as the dataset created in 2. For each ruling, we sample 5 answers to account for the probabilistic nature of the LLM and measure the consistency of GPT-4’s stance over repeated queries. An example of the answer’s format and content can be found in Appendix E.
We then use the embedding model trained in Section 3 to embed each of the arguments, and compute the cosine distance between GPT-4’s arguments and the other legal parties’. As previously mentioned, the embedding model was finetuned on the task of representing legal arguments so that opposing arguments will be further apart in space than concordant arguments. Due to this, we assume that this metric can be used to quantify the difference in alignment between different legal interpretations. Furthermore, since all of the text was generated by the same language model using the same data sources, we can mostly rule out the possibility that the embedding model is capturing stylistic differences rather than differences in legal intepretations. Figure 4 shows the results of this measurement, broken down by constitutional article. The next section discusses the results.
@@ 4 Results
Our evaluation in Section 2 demonstrates that GPT-4 performs adequately in the task of analyzing legal rulings. The model exhibits the capability to summarize and analyze text with a reasonable degree of completeness and accuracy, while showing a relatively low risk of hallucination. However, it is worth noting that providing additional background information on relevant jurisprudence could potentially enhance the quality of its output.
GPT-4 displays a notable proficiency in decoding legal reasoning presented by various parties involved in a case. It accurately discerns the different interpretations of cited articles and grasps the essence of the diverse positions advocated. Furthermore, the model demonstrates a high level of comprehension even when confronted with cases written in ”legalese.” This ability is particularly impressive given the complexity of specific legal domains, which typically require a certain degree of expertise to navigate effectively.
Despite these apparent capabilities, it remains challenging to determine whether the model is truly engaging in legal reasoning or if it is simply echoing the content it has been trained on. While GPT-4 demonstrates a high level of comprehension, its performance could be interpreted as a sophisticated form of paraphrasing rather than a deep understanding of constitutional principles and their nuanced interpretations. On the other hand, an alternative perspective may contend that the model has merely performed a basic legal task by outlining the main contents of a given text. This may not necessarily indicate a deep understanding of constitutional principles and their nuanced interpretations; it could be seen as simply paraphrasing the information it has been provided.
In any case, as demonstrated in previous research (Guha et al. 2024; Pont et al. 2023), our experiment corroborates that GPT-4 can be a valuable tool for accomplishing basic legal tasks, such as summarizing text and outlining relevant arguments. However, it is crucial to emphasize that these tasks should be performed under the supervision of legal experts.
@@ Appendix B Prompts
Extracting legal arguments from a case
@@
Asking GPT-4 to take a stance in a legal case